OpenStudy (anonymous):

Solve for x in x2 + 16x + 9 = 0

OpenStudy (johnweldon1993):

You've heard of the quadratic formula correct?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (luigi0210):

\[\frac{ -b+/- \sqrt{b^2-4ac} }{ 2a}\]

OpenStudy (johnweldon1993):

Yes as @Luigi0210 has shown ...What you have shown...that is the FORM of a quadratic The quadratic formula is the \[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (johnweldon1993):

You have an equation x2 + 16x + 9 = 0 With 3 coefficients 1x² + 16x + 9 1 is your FIRST coefficient...so assign that value to the letter 'a' 16 is your SECOND coefficient...so assign that value to the letter 'b' and 9 is your THIRD coefficient...so assign that to the letter 'c' So what will your equation look like?

OpenStudy (anonymous):

-16±√16^2-4(1)(9)/ 2(1) @johnweldon1993

OpenStudy (anonymous):

@Luigi0210 ?

OpenStudy (anonymous):

\[\frac{ -16 \pm \sqrt{16^{2}-4(1)(9)} }{ 2(1) }\] \[\frac{ -16 \pm \sqrt{256-36} }{ }\] \[\frac{ -16 \pm \sqrt{220} }{ 2 }\]

OpenStudy (espex):

Good job @CSmith17 , now what you will have is two equations, (-16/2)+ (radical/2) and (-16/2)-(radical/2), from that you will get your two roots for this problem.

OpenStudy (anonymous):

@RoseDryer is that final?

OpenStudy (johnweldon1993):

it is not final....you should equate those 2 equations that you will now have \[\frac{ -16 + \sqrt{220} }{ 2 }\] and \[\frac{ -16 - \sqrt{220} }{ 2 }\] solve those to get a number and those will be your answers to 'x'

OpenStudy (anonymous):

So i got -0.5 and -15.5 @johnweldon1993

OpenStudy (anonymous):

Use Quad.Equation

OpenStudy (johnweldon1993):

probably keep more decimal places and round only at the end... so it'd be - .058 and -15.4 that I get

OpenStudy (anonymous):

So -.058 and -15.4 would be the final answer? Or only one of them? @johnweldon1993

OpenStudy (johnweldon1993):

Both would be a correct answer for 'x' try it...try plugging both answers in for 'x' in your equation and you will get 0 (probably not EXACTLY because we DID round) but yes